A student buys a ticket to a concert at discount of 20%. His ticket cost $6.80.
What is full price?

Answer by likaaka(51)

Set your equation as so,
6.80/x=1-20%
6.80/x=80%, in other words 340 is 82% of what
Convert percentage to a decimal fraction
6.80/x=.80
6.80=.80x, solve for x
6.80/.80=x
x=8.50
Full price of the ticket is $8.50